# Molecular Quantum Computations on a Protein

### v1

@@ 1 Introduction
Simulations of the electronic structure of molecules are one of the most promising applications of quantum computing schleich2025cracking, weidman2024quantum, alexeev2024quantum, motta2022emerging. The recently embraced paradigm of quantum-centric computing, where classical high-performance computing (HPC) and quantum processing units (QPUs) work in concert with each other, showed that electronic structure simulations utilizing present day quantum computers can be scaled up to at least 36 molecular orbitals (MOs) and beyond 72 qubits robledo2024chemistry, shirakawa2025closedloopcalculationselectronicstructure bringing the possibility of useful quantum computing simulations closer to reality.
@@
Quantum-centric computing of the electronic structure of a molecular system is currently enabled using the quantum-selected configuration interaction (QSCI) method kanno2023quantum and its variant, the sample-based quantum diagonalization methodology (SQD) robledo2024chemistry, shirakawa2025closedloopcalculationselectronicstructure. The SQD methodology introduced an expansion of the QSCI method through an error mitigation procedure, called configuration recovery, that allows the recovery of samples of electron configurations that violate known symmetries of the ansatz state methodologyrobledo2024chemistry. The SQD methodology was further expanded with the introduction of the carryover procedure that preserves the most dominant electron configurations identified within each step of configuration recoveryshirakawa2025closedloopcalculationselectronicstructure.
@@
QSCI and SQD methodologies are both forms of selected configuration interaction (SCI) method and the hope is these methods will offer eventual quantum advantage through efficient sampling of relevant electron configurations. Despite a recent criticism of the sampling ability of the QSCI and SQD methodologiesreinholdt2025critical, the original SQD paper demonstrated through numerical experiments demonstrated the existence of quantum circuits enabling the electronic configuration sampling within the SQD method that outperforms the fully-classical heat-bath configuration interaction method (HCI)robledo2024chemistry. Moreover, a recent QSCI study demonstrated that the hardware runs are closely approaching the regime where quantum sampling can offer advantages over fully-classical simulationsweaving2025towards.
@@
Quantum-centric computing of electronic structure has already been shown to be promising in applications to studies of metal complexesrobledo2024chemistry, shirakawa2025closedloopcalculationselectronicstructure, nutzel2024solving, aromatic moleculessugisaki2025hamiltonian, piccinelli2025quantum, shirai2025enhancing, intermolecular interactionskaliakin2024accurate, sugisaki2025size, open-shell systemssugisaki2025hamiltonian, Ieva2024, excited states of molecular systems  shirai2025enhancing, barison2025quantum, systems of interest for material sciencebarroca2025surface, nogaki2025symmetry, wray2025convergence, electronic structure of molecules in implicit solventkaliakin2025implicit, quantum mechanics / molecular mechanics (QM/MM) bickley2025extending, bazayeva2025quantum, as well as free energy with molecular dynamics (MD) in explicit solvent bazayeva2025quantum. Moreover, the SQD method was evaluated against the W4-11 thermochemistry dataset that includes 124 total atomization, 83 bond dissociation, 20 isomerization, 505 heavy-atom transfer, and 13 nucleophilic substitution processes. This evaluation demonstrated the ability of the SQD method to describe a diverse set of chemical environments and reactions.raisuddin2025promise
@@
Application of quantum-centric electronic structure simulations to problems of relevance to healthcare and life sciences inevitably requires scaling of these simulations to biomolecular systems like proteinsbaiardi2023quantum, pal2024future, where applications of quantum mechanical (QM) methods were previously shown to provide high-accuracy in the prediction of disordered regions within a protein structureufimtsev2011charge as well as prediction of protein-ligand binding free energiesmobley2017predicting. However, despite substantial successes in scaling up QM simulations, existing classical computational resources are insufficient to carry out first-principles electronic structure simulations on protein systems by methods beyond the Hartree-Fock (HF), density functional theory (DFT), and Second-order Møller–Plesset perturbation theory (MP2)ching2021ultra, stocks2024breaking. Recent advances in GPU acceleration further pushed the scalability of coupled cluster singles, doubles, and perturbative triples (CCSD(T)) simulationsfajen2025accelerating, as well as full configuration interaction simulationsfales2020efficient, but these simulations on their own are still far from being scalable to large protein systems.
@@
Configuration interaction simulations are especially challenging to scale upgao2024distributed. In a 2013 study Yoshikawa et al.yoshikawa2013novel demonstrated the divide-and-conquer (DC) symmetry-adapted cluster configuration interaction (SACCI)yoshikawa2013divide simulation of photoactive yellow protein where only the chromophore and five surrounding amino-acid residues were treated with the SACCI method. The heat-bath configuration interaction (HCI) has been reported to be scaled up to only (12e, 190o), (14e, 108o) and (48e, 42o) active spacesholmes2016heat, sharma2017semistochastic, while the largest DMRG simulations are limited to (113e, 76o) and (63e, 58o).menczer2024parallel Finally, the advanced formulations of the complete active space self-consistent field (CASSCF) method including the many-body expanded full configuration interaction approachgreiner2024mbe and adaptive selected configuration approachlevine2020casscf are limited to simulations of (40e,42o), (50e, 50o), and (52e,52o) active spaces.
@@
Quantum mechanical calculations beyond the mean-field approach in large protein systems require the utilization of approximations enabling this class of large-scale simulations. Such approximations can be realized using fragment-based electronic structure methods. kitaura1999fragment, raghavachari2015accurate, erlanson2019fragment These algorithms are based on the decomposition of a prohibitively large system into manageable subsystems, and typically employ a quantum embedding scheme, where subsystems are treated with the higher-level method while lower-level methods are used to generate the initial mean-field object corresponding to the entire system sun2016quantum. Previous studies proposed the integration of quantum computing subroutines in the workflow of quantum embedding methods bauer2016hybrid, kreula2016few, bravyi2017complexity through feedback between subsystems and their environment based on the electronic density, the Green’s function, or the one-particle reduced density matrix. sun2016quantum
@@
Early success in integration of quantum computing with fragment-based methods was achieved with the variational quantum eigensolver (VQE) method peruzzo2014variational, motta2022emerging, bauer2020quantum in combination with the fragment molecular orbital (FMO), lim2024fragment divide and conquer (DC), yoshikawa2022quantum many-body expansion (MBE), ma2023multiscale density matrix embedding theory (DMET), rubin2016hybrid, kawashima2021optimizing, iijima2023towards, li2022toward, kirsopp2022quantum, cao2023ab, ma2023multiscale and other wavefunction-based embedding  vorwerk2022quantum, rossmannek2023quantum, gujarati2023quantum methods. However, utilization of VQE hindered the scalability of these simulations leading to these simulations being performed either with classical simulators of quantum circuits or being limited to quantum hardware execution using a maximum of 12 qubitsma2023multiscale.
@@
In our recent study we demonstrated that the SQD method can be combined with the DMET method allowing for scaling to simulations using 32 qubits, which substantially exceeds what was reported before with VQE-based methodologiesshajan2025toward. Moreover, unfragmented systems demonstrated in this study would require 41 and 89 qubits for a chain of 18 hydrogen atoms ($H_{18}$) and cyclohexane, respectively. Our initial success with the DMET-SQD approach lead to a recent publication by Patra et al. where researchers adapted the approach originally proposed in our study to simulate drug-like moleculespatra2025quantum. Moreover, Bierman et al. demonstrated another promising fragment-based simulation with SQD using quantum bootstrap embeddingbierman2025towards.
@@
In the present work, we show that embedding-based quantum centric simulations can be scaled even further to realistic protein models. To achieve such scalability, we moved beyond the original DMET formulation in favor of the wavefunction-based embedding method (EWF) which represents the further expansion of the base DMET technology. The EWF formulation offers three important benefits over the DMET formulation: 1) unlike the standard DMET method it allows for more flexible extension of the bath (surrounding entangled orbitals) through MP2-based calculations; 2) it eliminates the loop over the chemical potential allowing for a more efficient non-iterative procedure which further reduces the overall computational cost of simulationsnusspickel2022systematic; 3) because of the way EWF treats the fragment plus it’s bath, we can make the individual fragments as small as individual atoms, while still recovering high quality results. Figure 1 shows the EWF workflow adapted in present study and demonstrates how we integrate SQD and FCI solvers for quantum-centric calculations as well as how we utilize CCSD and MP2 solvers for classical benchmark
The Trp-cage miniprotein provides an ideal molecular benchmark for evaluating quantum embedding techniques due to its compact size, biological relevance, and the presence of multiple energetically distinct conformerssimmerling2002all. Although only  20 residues long ( 300 atoms), Trp-cage exhibits a well-defined hydrophobic core formed by the indole side chain of tryptophan, as well as a characteristic $\alpha$-helix and poly-proline loopmaruyama2023effect, ozono2021visualization. In this work, we focus on two representative structures: the lowest-energy folded conformer, and the unfolded formsimmerling2002all. Their differing geometries lead to distinct electronic environments, including variations in charge distribution, hydrogen bonding, and steric effects. Orbital localization and fragmentation of these geometries yield EWF clusters (fragment + expanded bath) spanning 6 to 33 molecular orbitals. Smaller fragments can be treated exactly, while larger ones exhibit substantial correlation and require more advanced solvers. This wide distribution of EWF cluster sizes makes Trp-cage a stringent test for the EWF methodology and highlights the need for a scalable solver strategy.
@@ 2.1 Embedded Wave Function (EWF)
The Embedded Wave Function (EWF) framework Booth2022, Booth2023 builds upon the Density Matrix Embedding Theory (DMET) formulation, which uses a mean-field calculation to define localized fragments and DMET bath orbitals that reproduce the one-particle entanglement between each fragment and its environment. EWF departs from traditional DMET in how correlation is treated and how global properties are reconstructed. Instead of enforcing self-consistency through matching one-particle density matrices, each fragment-and-bath EWF cluster is solved at a correlated level, and the resulting one- and two-body density matrices are projected back into the full molecular orbital basis. These projected quantities are then combined across EWF clusters to obtain global expectation values, including the total energy.
@@ 2.1.2 Interacting Bath Expansion
In the DMET literature chen2014intermediate, ma2023multiscale, rubin2016hybrid, the fragment and its entangled surroundings are commonly referred to as the impurity. In the EWF framework, the initial DMET bath is expanded in a manner analogous to pair natural orbitals, introducing additional occupied and virtual states that systematically recover the local excitations needed for an accurate correlated fragment description and an improved representation of fragment–environment entanglement. To identify the most relevant additions to this space, the local interacting subspace from second-order Møller–Plesset perturbation theory (MP2) is costructed. Two subspace MP2 calculations are performed for each fragment: one in which the occupied excitations are restricted to the DMET impurity space, and another in which the virtual excitations are similarly restricted. The resulting correlated one- and two-body density matrices are used to form bath natural orbitals spanning the corresponding occupied and virtual environment spaces. Their occupation numbers quantify the strength of correlation-driven coupling to the fragment. Natural orbitals with occupations larger than a threshold $\eta$ (for virtual) or smaller than $2-\eta$ (for occupied) are retained and included with the fragment and original DMET bath to define the final correlated EWF impurity. Following this expansion, the combined fragment–bath unit is referred to as an EWF cluster, consistent with the original EWF terminology Booth2022.
@@ 2.2 Sample-based quantum diagonalization (SQD) and extended SQD (ext-SQD)
Recent developments in hybrid quantum-classical algorithms have unlocked new possibilities for scaling up applications of quantum computing for chemistry. These are based on the classical processing of individual quantum samples, such as the quantum-selected configuration interaction method kanno2023quantum, which extended ideas from the classical configuration interaction methods to the quantum domain, and the sample-based quantum diagonalization (SQD) workflow, which allowed scaling for up to 85 qubits robledo2024chemistry, yu2025quantum for molecular and fermionic lattice models. We sample candidate electron configurations, to be used within the SQD workflow, from the local unitary cluster Jastrow (LUCJ) motta2023bridging ansatz. Such an approach leverages the idea that appropriate quantum circuits can in principle produce an electron configuration subspace that is significantly more compact than a subspace produced with a fully-classical heuristic, while still predicting the accurate total energy of the target system. The sampling of electron configurations with the LUCJ circuit was demonstrated through numerical experiments in the original SQD paper. robledo2024chemistry The quantum-centric aspect of SQD involves using a quantum computer to generate candidate configurations, and classical high-performance computing (HPC) resources to carry out error mitigation at the level of individual samples and to perform diagonalization of a projected Hamiltonian.
After identifying a subspace with SQD, we further improve accuracy by first identifying the most relevant configurations within the obtained subspaces and then extend these configurations through single excitation operators (so-called extended-SQD barison2025quantum). The resulting subspace allows for even more extensive capture of relevant configurations and improves the quality of the resulting total energy and reduced density matrices.
@@ 3.1.1 Conformer definition
In this work, we calculate the relative energy between the most energetically stable and most energetically unstable conformers of Trp-cage (henceforth called the “folded” and “unfolded” conformers), as defined in a previous study by Simmerling et al.Simmerling2002 We generated the protonated structures of these two systems using the H++ anandakrishnan2012h++, myers2006simple, gordon33h++ web server at pH = 7. During the protonation process we only add hydrogen atoms that correspond to protonation of amino acids at a given pH. However, we are not using explicit or implicit solvents, and the simulations are performed in the gas phase. Upon protonation, the overall charge of the system is +1. To date, no DMET or EWF studies have demonstrated the simulation of charged species Booth2022, Booth2023. We believe that the charged cases represent a challenge for these methods and require further detailed study. To maintain the most biologically relevant protonation while changing the nature of the overall system to a zwitterionic form (where individual charged residues lead to an overall neutral charge of the system), we doprotonated LYS8 via removal of one of the hydrogen atoms from the NH3 group, for both the folded and unfolded conformers. We chose LYS8 for modification because it is located on the surface of the protein in both conformers and it is not involved in critical non-covalent interactions. We treated the system as a closed-shell singlet spin state.

### v2

@@ 1 Introduction
Simulations of the electronic structure of molecules are one of the most promising applications of quantum computing 74, 94, 2, 50. The recently embraced paradigm of quantum-centric computing, where classical high-performance computing (HPC) and quantum processing units (QPUs) work in concert with each other, showed that electronic structure simulations utilizing present day quantum computers can be scaled up to at least 36 molecular orbitals (MOs) and beyond 72 qubits 70, 78 bringing the possibility of useful quantum computing simulations closer to reality.
@@
Quantum-centric computing of the electronic structure of a molecular system is currently enabled using the quantum-selected configuration interaction (QSCI) method 32 and its variant, the sample-based quantum diagonalization methodology (SQD) 70, 78. The SQD methodology introduced an expansion of the QSCI method through an error mitigation procedure, called configuration recovery, that allows the recovery of samples of electron configurations that violate known symmetries of the ansatz state methodology70. The SQD methodology was further expanded with the introduction of the carryover procedure that preserves the most dominant electron configurations identified within each step of configuration recovery78.
@@
QSCI and SQD methodologies are both forms of selected configuration interaction (SCI) method and the hope is these methods will offer eventual quantum advantage through efficient sampling of relevant electron configurations. Despite a recent criticism of the sampling ability of the QSCI and SQD methodologies68, the original SQD paper demonstrated through numerical experiments demonstrated the existence of quantum circuits enabling the electronic configuration sampling within the SQD method that outperforms the fully-classical heat-bath configuration interaction method (HCI)70. Moreover, a recent QSCI study demonstrated that the hardware runs are closely approaching the regime where quantum sampling can offer advantages over fully-classical simulations93.
@@
Quantum-centric computing of electronic structure has already been shown to be promising in applications to studies of metal complexes70, 78, 58, aromatic molecules84, 64, 77, intermolecular interactions31, 85, open-shell systems84, 40, excited states of molecular systems  77, 5, systems of interest for material science6, 54, 96, electronic structure of molecules in implicit solvent30, quantum mechanics / molecular mechanics (QM/MM) 10, 9, as well as free energy with molecular dynamics (MD) in explicit solvent 9. Moreover, the SQD method was evaluated against the W4-11 thermochemistry dataset that includes 124 total atomization, 83 bond dissociation, 20 isomerization, 505 heavy-atom transfer, and 13 nucleophilic substitution processes. This evaluation demonstrated the ability of the SQD method to describe a diverse set of chemical environments and reactions.67
@@
Application of quantum-centric electronic structure simulations to problems of relevance to healthcare and life sciences inevitably requires scaling of these simulations to biomolecular systems like proteins4, 61, where applications of quantum mechanical (QM) methods were previously shown to provide high-accuracy in the prediction of disordered regions within a protein structure90 as well as prediction of protein-ligand binding free energies48. However, despite substantial successes in scaling up QM simulations, existing classical computational resources are insufficient to carry out first-principles electronic structure simulations on protein systems by methods beyond the Hartree-Fock (HF), density functional theory (DFT), and Second-order Møller–Plesset perturbation theory (MP2)15, 83. Recent advances in GPU acceleration further pushed the scalability of coupled cluster singles, doubles, and perturbative triples (CCSD(T)) simulations19, as well as full configuration interaction simulations20, but these simulations on their own are still far from being scalable to large protein systems.
@@
Configuration interaction simulations are especially challenging to scale up22. In a 2013 study Yoshikawa et al.98 demonstrated the divide-and-conquer (DC) symmetry-adapted cluster configuration interaction (SACCI)99 simulation of photoactive yellow protein where only the chromophore and five surrounding amino-acid residues were treated with the SACCI method. The heat-bath configuration interaction (HCI) has been reported to be scaled up to only (12e, 190o), (14e, 108o) and (48e, 42o) active spaces26, 76, while the largest DMRG simulations are limited to (113e, 76o) and (63e, 58o).47 Finally, the advanced formulations of the complete active space self-consistent field (CASSCF) method including the many-body expanded full configuration interaction approach24 and adaptive selected configuration approach37 are limited to simulations of (40e,42o), (50e, 50o), and (52e,52o) active spaces.
@@
Quantum mechanical calculations beyond the mean-field approach in large protein systems require the utilization of approximations enabling this class of large-scale simulations. Such approximations can be realized using fragment-based electronic structure methods. 35, 66, 18 These algorithms are based on the decomposition of a prohibitively large system into manageable subsystems, and typically employ a quantum embedding scheme, where subsystems are treated with the higher-level method while lower-level methods are used to generate the initial mean-field object corresponding to the entire system 89. Previous studies proposed the integration of quantum computing subroutines in the workflow of quantum embedding methods 8, 36, 12 through feedback between subsystems and their environment based on the electronic density, the Green’s function, or the one-particle reduced density matrix. 89
@@
Early success in integration of quantum computing with fragment-based methods was achieved with the variational quantum eigensolver (VQE) method 63, 50, 7 in combination with the fragment molecular orbital (FMO), 41 divide and conquer (DC), 100 many-body expansion (MBE), 44 density matrix embedding theory (DMET), 72, 33, 28, 39, 34, 13, 44 and other wavefunction-based embedding  91, 71, 25 methods. However, utilization of VQE hindered the scalability of these simulations leading to these simulations being performed either with classical simulators of quantum circuits or being limited to quantum hardware execution using a maximum of 12 qubits44.
@@
In our recent study we demonstrated that the SQD method can be combined with the DMET method allowing for scaling to simulations using 32 qubits, which substantially exceeds what was reported before with VQE-based methodologies75. Moreover, unfragmented systems demonstrated in this study would require 41 and 89 qubits for a chain of 18 hydrogen atoms ($H_{18}$) and cyclohexane, respectively. Our initial success with the DMET-SQD approach lead to a recent publication by Patra et al. where researchers adapted the approach originally proposed in our study to simulate drug-like molecules62. Moreover, Bierman et al. demonstrated another promising fragment-based simulation with SQD using quantum bootstrap embedding11.
@@
In the present work, we show that embedding-based quantum centric simulations can be scaled even further to realistic protein models. To achieve such scalability, we moved beyond the original DMET formulation in favor of the wavefunction-based embedding method (EWF) which represents the further expansion of the base DMET technology. The EWF formulation offers three important benefits over the DMET formulation: 1) unlike the standard DMET method it allows for more flexible extension of the bath (surrounding entangled orbitals) through MP2-based calculations; 2) it eliminates the loop over the chemical potential allowing for a more efficient non-iterative procedure which further reduces the overall computational cost of simulations56; 3) because of the way EWF treats the fragment plus it’s bath, we can make the individual fragments as small as individual atoms, while still recovering high quality results. Figure 1 shows the EWF workflow adapted in present study and demonstrates how we integrate SQD and FCI solvers for quantum-centric calculations as well as how we utilize CCSD and MP2 solvers for classical benchmark
The Trp-cage miniprotein provides an ideal molecular benchmark for evaluating quantum embedding techniques due to its compact size, biological relevance, and the presence of multiple energetically distinct conformers82. Although only  20 residues long ( 300 atoms), Trp-cage exhibits a well-defined hydrophobic core formed by the indole side chain of tryptophan, as well as a characteristic $\alpha$-helix and poly-proline loop46, 60. In this work, we focus on two representative structures: the lowest-energy folded conformer, and the unfolded form82. Their differing geometries lead to distinct electronic environments, including variations in charge distribution, hydrogen bonding, and steric effects. Orbital localization and fragmentation of these geometries yield EWF clusters (fragment + expanded bath) spanning 6 to 33 molecular orbitals. Smaller fragments can be treated exactly, while larger ones exhibit substantial correlation and require more advanced solvers. This wide distribution of EWF cluster sizes makes Trp-cage a stringent test for the EWF methodology and highlights the need for a scalable solver strategy.
@@ 2.1 Embedded Wave Function (EWF)
The Embedded Wave Function (EWF) framework 55, 57 builds upon the Density Matrix Embedding Theory (DMET) formulation, which uses a mean-field calculation to define localized fragments and DMET bath orbitals that reproduce the one-particle entanglement between each fragment and its environment. EWF departs from traditional DMET in how correlation is treated and how global properties are reconstructed. Instead of enforcing self-consistency through matching one-particle density matrices, each fragment-and-bath EWF cluster is solved at a correlated level, and the resulting one- and two-body density matrices are projected back into the full molecular orbital basis. These projected quantities are then combined across EWF clusters to obtain global expectation values, including the total energy.
@@ 2.1.2 Interacting Bath Expansion
In the DMET literature 14, 44, 72, the fragment and its entangled surroundings are commonly referred to as the impurity. In the EWF framework, the initial DMET bath is expanded in a manner analogous to pair natural orbitals, introducing additional occupied and virtual states that systematically recover the local excitations needed for an accurate correlated fragment description and an improved representation of fragment–environment entanglement. To identify the most relevant additions to this space, the local interacting subspace from second-order Møller–Plesset perturbation theory (MP2) is costructed. Two subspace MP2 calculations are performed for each fragment: one in which the occupied excitations are restricted to the DMET impurity space, and another in which the virtual excitations are similarly restricted. The resulting correlated one- and two-body density matrices are used to form bath natural orbitals spanning the corresponding occupied and virtual environment spaces. Their occupation numbers quantify the strength of correlation-driven coupling to the fragment. Natural orbitals with occupations larger than a threshold $\eta$ (for virtual) or smaller than $2-\eta$ (for occupied) are retained and included with the fragment and original DMET bath to define the final correlated EWF impurity. Following this expansion, the combined fragment–bath unit is referred to as an EWF cluster, consistent with the original EWF terminology 55.
@@ 2.2 Sample-based quantum diagonalization (SQD) and extended SQD (ext-SQD)
Recent developments in hybrid quantum-classical algorithms have unlocked new possibilities for scaling up applications of quantum computing for chemistry. These are based on the classical processing of individual quantum samples, such as the quantum-selected configuration interaction method 32, which extended ideas from the classical configuration interaction methods to the quantum domain, and the sample-based quantum diagonalization (SQD) workflow, which allowed scaling for up to 85 qubits 70, 101 for molecular and fermionic lattice models. We sample candidate electron configurations, to be used within the SQD workflow, from the local unitary cluster Jastrow (LUCJ) 49 ansatz. Such an approach leverages the idea that appropriate quantum circuits can in principle produce an electron configuration subspace that is significantly more compact than a subspace produced with a fully-classical heuristic, while still predicting the accurate total energy of the target system. The sampling of electron configurations with the LUCJ circuit was demonstrated through numerical experiments in the original SQD paper. 70 The quantum-centric aspect of SQD involves using a quantum computer to generate candidate configurations, and classical high-performance computing (HPC) resources to carry out error mitigation at the level of individual samples and to perform diagonalization of a projected Hamiltonian.
After identifying a subspace with SQD, we further improve accuracy by first identifying the most relevant configurations within the obtained subspaces and then extend these configurations through single excitation operators (so-called extended-SQD 5). The resulting subspace allows for even more extensive capture of relevant configurations and improves the quality of the resulting total energy and reduced density matrices.
@@ 3.1.1 Conformer definition
In this work, we calculate the relative energy between the most energetically stable and most energetically unstable conformers of Trp-cage (henceforth called the “folded” and “unfolded” conformers), as defined in a previous study by Simmerling et al.81 We generated the protonated structures of these two systems using the H++ 3, 52, 23 web server at pH = 7. During the protonation process we only add hydrogen atoms that correspond to protonation of amino acids at a given pH. However, we are not using explicit or implicit solvents, and the simulations are performed in the gas phase. Upon protonation, the overall charge of the system is +1. To date, no DMET or EWF studies have demonstrated the simulation of charged species 55, 57. We believe that the charged cases represent a challenge for these methods and require further detailed study. To maintain the most biologically relevant protonation while changing the nature of the overall system to a zwitterionic form (where individual charged residues lead to an overall neutral charge of the system), we doprotonated LYS8 via removal of one of the hydrogen atoms from the NH3 group, for both the folded and unfolded conformers. We chose LYS8 for modification because it is located on the surface of the protein in both conformers and it is not involved in critical non-covalent interactions. We treated the system as a closed-shell singlet spin state.